OpenStudy (anonymous):

10 ÷ (-5) x -2 x -8 = ____

OpenStudy (smartnerd1111):

10/−5−(2)(−8) =−2−(2)(−8) =−2−(−16) =14

OpenStudy (anonymous):

k thx

OpenStudy (anonymous):

It said that it isn't right @smartnerd1111

OpenStudy (anonymous):

It said it was -32

OpenStudy (smartnerd1111):

oh